# Phantom scalar field with arbitrary potential: accelerating scaling attractors

### v1

@@ Abstract
In this article, we investigate the dynamics of a phantom scalar field with an arbitrary potential, focusing on accelerating scaling solutions of cosmological relevance. We consider both uncoupled and coupled cosmological scenarios. In the latter case, the coupling between phantom dark energy and dark matter is motivated by the warm inflationary paradigm, with the dissipation coefficient assumed to be either constant or variable. The evolution equations of our coupled and uncoupled cosmological models are written in the form of autonomous systems, whose stability is studied using methods of qualitative analysis of dynamical systems. For this analysis, the only requirement imposed on the phantom scalar-field potential is that a specific dynamical variable, defined in terms of the potential and its derivative, must be invertible. We show that the uncoupled phantom cosmological model cannot accommodate any accelerated scaling solution, while such solutions do exist in the coupled scenario, for both constant and variable dissipation coefficients. Although there is a limitation to these scaling solutions - specifically, the current stage of accelerated expansion is not preceded by a long enough matter-dominated era - our results show that the existence of a direct coupling between phantom dark energy and dark matter yields great potential for addressing the cosmic coincidence problem.
@@ I Introduction
In the present article, we particularly focus on the phantom scalar-field model [15, 16], with the underlying gravitational field solely described by GR. Since the total energy of the phantom field is unbounded from below, this model is appropriate for describing only the late-time evolution of the Universe.
@@
An interacting cosmological scenario (see, for instance, Refs. [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 50, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97]) is a generalized cosmological framework offering many interesting outcomes. From a rational point of view, there is no objection to considering interactive scenarios, as long as they are not ruled out by observations. The aftermath of a direct interaction between the dark sectors is far-reaching — such an interaction can alleviate the cosmic coincidence problem [19, 21, 26, 23] and the Hubble constant tension [98, 42, 99, 53, 55].
@@
One of the main foci of the present article is to search for late-time accelerating scaling attractors in phantom scalar field models, since, according to the existing literature in this direction, only a few articles have found such a possibility [101, 102, 103]. Our work presents a novel contribution to the literature of phantom cosmologies by allowing the potential $V(\phi)$ for the phantom scalar field to be quite general, both for the uncoupled and coupled scenarios. The only requirement on this potential is that a specific dynamical variable $\lambda(\phi)$, which is defined in terms of $V(\phi)$ and its derivative $\partial V/\partial\phi$ (see Eq. (9)), must be invertible. For such a rather generic potential, our findings indicate that in the uncoupled phantom model, no accelerating scaling attractors can exist. In contrast, such solutions are present in the coupled phantom scenario. This is an interesting result, suggesting that a direct, non-gravitational interaction between the dark components of the Universe might help to explain recent observational evidence.
@@ IV Coupled phantom dark energy
Let us now move on to the coupled case, assuming the interaction term to be of the dissipative type [108]
@@
where $\Gamma$ is the dissipation coefficient.
@@ IV.1 Constant dissipation coefficient
For such choice of $\Gamma$, the evolution equations (2)–(5) cannot be written as a three-dimensional dynamical system, as in the uncoupled case (see Eqs. (10)), since the interaction term $Q$ between DE and DM cannot be expressed as a function of the variables $x$, $y$, and $\lambda$, defined in Eqs. (7) and (9); an additional variable $z$ is required to close the dynamical system, which therefore becomes four-dimensional.
@@
Let us start with the critical surface $C_{3}$. We consider a point $C_{3}(0,y_{c},1,\lambda_{c})$ on this surface, where $0\leq y_{c}\leq 1$ and $\lambda_{c}\in\mathbb{R}$. At this point, the Jacobian matrix of the dynamical system (23) has the eigenvalues $\Lambda_{1}=\alpha(y_{c}^{2}-1)$ and $\Lambda_{2,3,4}=0$. Since more than two eigenvalues are zero, the linear theory does not suffice to assess the stability of this critical point, and consequently, we have to resort to alternative methods. [^4] If only two eigenvalues were zero (those whose corresponding eigenvectors are tangent to the surface $C_{3}$), this surface would be normally hyperbolic, and linear theory would suffice to study its stability. In this scenario, the signs of the two nonzero eigenvalues would determine whether $C_{3}$ is an attractor, saddle, or repellor in the remaining directions. In Appendix A, using the center manifold theory, we establish that this critical point is not an attractor for $0\leq y_{c}<1$, $\lambda_{c}\in\mathbb{R}$, and any value of $\alpha$. For $y_{c}=1$ and $\lambda_{c}\in\mathbb{R}$, on the contrary, this critical point is indeed an attractor for $\alpha>0$, but, in this case, $\Omega_{\phi}=1$ and $\Omega_{\texttt{DM}}=0$ (see Table 2); therefore, it does not correspond to a scaling solution, but rather to a final state entirely dominated by phantom DE.
@@ IV.2 Variable dissipation coefficient
As in the case of constant $\Gamma$, the interaction term $Q$ between DE and DM cannot be expressed as a function of the variables $x$, $y$, and $\lambda$, defined in Eqs. (7) and (9). An additional variable $z$ is needed to close the dynamical system, which we choose to be given again by Eq. (21). Recall that such a choice, which compactifies the phase space in the $z$ direction, leads to a divergent interaction term in the evolution equation for $x$ (see discussion following Eq. (21)). This divergence is removed by choosing a new “time” variable $\sigma$, given by
@@
Linear theory is not enough to assess the stability of $E_{3}$, since the Jacobian matrix has more than two zero eigenvalues. Therefore, we resort to the center manifold theory to establish that this critical surface is an attractor only along the line $y=1$, $\lambda\in\mathbb{R}$ (see Appendix B), implying that the corresponding final state is entirely dominated by phantom dark energy.
Regarding the critical line $E_{6}$, its stability properties can be established using linear theory alone (see Appendix B), since only one of the eigenvalues of the Jacobian matrix is zero (the one whose corresponding eigenvector is tangent to the line $E_{6}$). Two other eigenvalues have negative real parts, while the fourth is negative if $x_{c}f^{\prime}(0)>0$, where $x_{c}$ is the $x$-coordinate of a specific point on the line $E_{6}$ and the prime denotes the derivative with respect to $\lambda$. Consequently, we conclude that the critical line $E_{6}$ is an attractor if the above condition is satisfied. However, similarly to the case of the coupled model with a constant dissipation coefficient (see Fig. 1 and discussion around it), phase-space trajectories that end at $E_{6}$ in points with an $x$-coordinate lying in the interval $0\ll|x|<1/\sqrt{2}$ (i.e., points corresponding to accelerated scaling solutions) do not pass close to the critical line $E_{1}$, which corresponds to a matter-dominated solution.
The situation here is similar to that found in the coupled model with a constant dissipation coefficient. We indeed find a final stage of evolution corresponding to a scaling solution. Still, this stage is not preceded by a long enough matter-dominated era, as required by cosmological observations. However, despite this limitation, the appearance of scaling solutions in our coupled phantom DE cosmological model is certainly appealing. Finding such solutions in phantom models has proven to be really challenging, with only a few examples available for specific potentials and interaction terms [101, 102, 103]. Therefore, the results of the present study, achieved for a generic potential, will certainly motivate future research in this direction.
@@ V Conclusions
Acknowledgements.
SH was supported by a senior research fellowship from the University Grants Commission (UGC), Govt. of India (NTA Ref. No: 201610019097). SP and TS were partially supported by the Department of Science and Technology (DST), Govt. of India under the Scheme “Fund for Improvement of S&T Infrastructure (FIST)” (File No. SR/FST/MS-I/2019/41). PS acknowledges support from CIDMA–Center for Research and Development in Mathematics and Applications under the Portuguese Foundation for Science and Technology Multi-Annual Financing Program for R&D Units, grants UID/4106/2025 and UID/PRR/4106/2025.

### v2

# Phantom cosmology with arbitrary potential: New accelerating scaling attractors

@@ Abstract
In this article, we investigate the existence of accelerating scaling solutions in coupled phantom cosmology without assuming any specific potential for the phantom scalar field. The coupling between phantom dark energy and dark matter is motivated by the warm inflationary paradigm, with the dissipation coefficient assumed to be either constant or variable. The evolution equations are written in the form of autonomous systems, whose stability is studied using methods of qualitative analysis of dynamical systems. For this analysis, the only requirement imposed on the otherwise arbitrary phantom potential is that a particular dynamical variable, defined in terms of the potential and its derivative, must be invertible. For such a generic potential, we show that accelerated scaling solutions do exist, for both constant and variable dissipation coefficients. Although there is a limitation to these scaling solutions – specifically, the current stage of accelerated expansion is not preceded by a long enough matter-dominated era – our results show that the existence of a direct coupling between phantom dark energy and dark matter yields great potential for addressing the cosmic coincidence problem.
@@ I Introduction
In the present article, we particularly focus on the phantom scalar-field model [15, 16, 17], with the underlying gravitational field solely described by GR. Since the total energy of the phantom field is unbounded from below, this model is appropriate for describing only the late-time evolution of the Universe.
@@
An interacting cosmological scenario (see, for instance, Refs. [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86]) is a generalized cosmological framework offering many interesting outcomes. From a rational point of view, there is no objection to considering interactive scenarios, as long as they are not ruled out by observations. The aftermath of a direct interaction between the dark sectors is far-reaching — such an interaction can alleviate the cosmic coincidence problem [19, 20, 25] and the Hubble constant tension [87, 36, 88, 45].
@@
One of the main foci of the present article is to search for late-time accelerating scaling attractors in phantom scalar field models, since, according to the existing literature in this direction, only a few articles have found such a possibility [90, 91, 92, 93]. Our work presents a novel contribution to the literature of phantom cosmologies by allowing the potential $V(\phi)$ for the phantom scalar field to be quite general, both for the uncoupled and coupled scenarios. The only requirement on this potential is that a specific dynamical variable $\lambda(\phi)$, which is defined in terms of $V(\phi)$ and its derivative $\partial V/\partial\phi$ (see Eq. (9)), must be invertible. For such a rather generic potential, our findings indicate that in the uncoupled phantom model, no accelerating scaling attractors can exist. In contrast, such solutions are present in the coupled phantom scenario. This is an interesting result, suggesting that a direct, non-gravitational interaction between the dark components of the Universe might help to explain recent observational evidence.
@@ IV Coupled phantom dark energy
Let us now move on to the coupled case. Expressed in terms of the dimensionless variables $x$, $y$, and $\lambda$, introduced in the previous section, the evolution equations (2)-(5) read
$x_{N}$ $=-\frac{\sqrt{6}}{2}\lambda y^{2}-\frac{3}{2}x(1+x^{2}+y^{2})-\frac{\kappa^{2}Q}{6H^{3}x},$ (19a)
$\lambda_{N}$ $=-\sqrt{6}f(\lambda)x,$ (19c)
where $f(\lambda)$ is defined in Eq. (11). Depending on the functional form of $Q$, it may be necessary to introduce an extra dimensionless variable to render the system autonomous.
We assume the interaction term $Q$ to be of the dissipative type [98]
@@
where $\Gamma$ is the dissipation coefficient. It has the dimension of mass, implying that $Q$ has the dimension of (mass)5, which is consistent with the dimensions of the terms $\dot{\rho}_{\texttt{DM}}$ ($\dot{\rho}_{\phi}$) and $3H\rho_{\texttt{DM}}$ ($3H\rho_{\phi}$) in the conservation equations, both with dimension (mass)5.
@@ IV.1 Constant dissipation coefficient
For such choice of $\Gamma$, the evolution equations (2)–(5) cannot be written as a three-dimensional dynamical system, as in the uncoupled case (see Eqs. (10)), since the interaction term $Q$ between DE and DM cannot be expressed as a function of the variables $x$, $y$, and $\lambda$, defined in Eqs. (7) and (9); an additional variable $z$ is required to close the dynamical system (19), which therefore becomes four-dimensional.
@@
It is worth mentioning that this choice of the dimensionless variable $z$ also guarantees that the hypersurface $\Omega_{\texttt{DM}}=1+x^{2}-y^{2}=0$ is an invariant manifold of the four-dimensional dynamical system, meaning that no phase-space trajectory can cross this hypersurface and enter the unphysical region of negative values of $\Omega_{\texttt{DM}}$.
@@
Let us start with the critical surface $C_{3}$. We consider a point $C_{3}(0,y_{c},1,\lambda_{c})$ on this surface, where $0\leq y_{c}\leq 1$ and $\lambda_{c}\in\mathbb{R}$. At this point, the Jacobian matrix of the dynamical system (24) has the eigenvalues $\Lambda_{1}=\alpha(y_{c}^{2}-1)$ and $\Lambda_{2,3,4}=0$. Since more than two eigenvalues are zero, the linear theory does not suffice to assess the stability of this critical point, and consequently, we have to resort to alternative methods. [^4] If only two eigenvalues were zero (those whose corresponding eigenvectors are tangent to the surface $C_{3}$), this surface would be normally hyperbolic, and linear theory would suffice to study its stability. In this scenario, the signs of the two nonzero eigenvalues would determine whether $C_{3}$ is an attractor, saddle, or repellor in the remaining directions. In A, using the center manifold theory, we establish that this critical point is not an attractor for $0\leq y_{c}<1$, $\lambda_{c}\in\mathbb{R}$, and any value of $\alpha$. For $y_{c}=1$ and $\lambda_{c}\in\mathbb{R}$, on the contrary, this critical point is indeed an attractor for $\alpha>0$, but, in this case, $\Omega_{\phi}=1$ and $\Omega_{\texttt{DM}}=0$ (see Table 2); therefore, it does not correspond to a scaling solution, but rather to a final state entirely dominated by phantom DE.
@@
The phase space associated with the dynamical system (24) is not bounded in the $x$, $y$, and $\lambda$ directions. Consequently, the existence of critical points at infinity cannot be ruled out. The asymptotic structure of the phase space and the stability of these putative points are investigated in C. We find out that there are two critical points at infinity, both exhibiting unstable behavior. Consequently, the interaction term $Q$, given by Eq. (20), with a constant dissipation coefficient effectively avoids Big-Rip singularities.
@@ IV.2 Variable dissipation coefficient
As in the case of constant $\Gamma$, the interaction term $Q$ between DE and DM cannot be expressed as a function of the variables $x$, $y$, and $\lambda$, defined in Eqs. (7) and (9). An additional variable $z$ is needed to close the dynamical system (19), which we choose to be given again by Eq. (22). Recall that such a choice, which compactifies the phase space in the $z$ direction, leads to a divergent interaction term in the evolution equation for $x$ (see discussion following Eq. (22)). This divergence is removed by choosing a new “time” variable $\sigma$, given by
@@
Linear theory is not enough to assess the stability of $E_{3}$, since the Jacobian matrix has more than two zero eigenvalues. Therefore, we resort to the center manifold theory to establish that this critical surface is an attractor only along the line $y=1$, $\lambda\in\mathbb{R}$ (see B), implying that the corresponding final state is entirely dominated by phantom dark energy.
Regarding the critical line $E_{6}$, its stability properties can be established using linear theory alone (see B), since only one of the eigenvalues of the Jacobian matrix is zero (the one whose corresponding eigenvector is tangent to the line $E_{6}$). Two other eigenvalues have negative real parts, while the fourth is negative if $x_{c}f^{\prime}(0)>0$, where $x_{c}$ is the $x$-coordinate of a specific point on the line $E_{6}$ and the prime denotes the derivative with respect to $\lambda$. Consequently, we conclude that the critical line $E_{6}$ is an attractor if the above condition is satisfied. However, similarly to the case of the coupled model with a constant dissipation coefficient (see Fig. 1 and discussion around it), phase-space trajectories that end at $E_{6}$ in points with an $x$-coordinate lying in the interval $0\ll|x|<1/\sqrt{2}$ (i.e., points corresponding to accelerated scaling solutions) do not pass close to the critical line $E_{1}$, which corresponds to a matter-dominated solution.
The situation here is similar to that found in the coupled model with a constant dissipation coefficient. We indeed find a final stage of evolution corresponding to a scaling solution. Still, this stage is not preceded by a long enough matter-dominated era, as required by cosmological observations. However, despite this limitation, the appearance of scaling solutions in our coupled phantom DE cosmological model is certainly appealing. Finding such solutions in phantom models has proven to be really challenging, with only a few examples available for specific potentials and interaction terms [90, 91, 92, 93]. Therefore, the results of the present study, achieved for a generic potential, will certainly motivate future research in this direction.
The asymptotic structure of the phase space and the stability of the critical points located at infinity are investigated in C. As for the case of a constant dissipative coefficient, we find that the interaction term $Q$, given by Eq. (20), with a variable dissipation coefficient effectively avoids Big-Rip singularities.
@@ V Conclusions
Acknowledgements
We thank the anonymous referee for useful comments and suggestions. SH was supported by a senior research fellowship from the University Grants Commission (UGC), Govt. of India (NTA Ref. No: 201610019097). SP and TS were partially supported by the Department of Science and Technology (DST), Govt. of India under the Scheme “Fund for Improvement of S&T Infrastructure (FIST)” (File No. SR/FST/MS-I/2019/41). PS acknowledges support from Center for Research and Development in Mathematics and Applications (CIDMA) under the Portuguese Foundation for Science and Technology grants UID/04106/2025 (https://doi.org/10.54499/UID/04106/2025) and UID/PRR/04106/2025 (https://doi.org/10.54499/UID/PRR/04106/2025).
@@ Appendix B Stability of the critical surface $E_{3}$ and the critical line $E_{6}$
Appendix C Stability of the critical points lying at infinity
In this Appendix, we investigate the asymptotic structure of the phase space and the stability of critical points located at infinity, both for the case of a constant dissipation coefficient and for the case of a variable coefficient. To that end we employ the Poincaré compactification technique [108, 109, 110, 89, 92].
Let us start with the case of a constant dissipation coefficient.
For convenience, we express the dynamical system (24) in terms of its associated vector field $F=(P,Q,R,S)$, where the individual components are given by the right-hand side of Eqs. (24a)–(24d), respectively. Motivated by the dynamical requirements and without affecting the results obtained in the previous sections, we choose $f(\lambda)$ as a polynomial admitting both a zero and a non-zero root. Accordingly, we consider $f(\lambda)=\delta\lambda(\lambda-\overline{\lambda})$, with non-zero real constants $\delta$ and $\overline{\lambda}$, a form [^7] We choose $f(\lambda)$ to be a second-degree polynomial for simplicity and concreteness. Nevertheless, $f(\lambda)$ itself is not restricted to this specific form and may, in principle, be any function possessing at least one zero and one non-zero root. However, the present analysis is carried out only for the polynomial case. supported by a wide class of scalar-field potentials studied in the literature [89].
To analyze the asymptotic dynamics, we first map the vector field $F$ onto the upper four-dimensional sphere $\mathbb{S}^{4+}$ through a central projection. This is achieved by introducing the projective coordinates $(x_{1},x_{2},x_{3},x_{4},x_{5})$ defined as $x_{1}=\frac{x}{\Delta}$, $x_{2}=\frac{y}{\Delta}$, $x_{3}=\frac{z}{\Delta}$, $x_{4}=\frac{\lambda}{\Delta}$, $x_{5}=\frac{1}{\Delta}$, where $\Delta=\sqrt{1+x^{2}+y^{2}+z^{2}+\lambda^{2}}$. The stability properties are then examined in five local coordinate charts $(V_{i},\psi_{i})$, $i=1,\dots,5$, with $V_{i}=\left\{(x_{1},x_{2},x_{3},x_{4},x_{5})\in\mathbb{S}^{4}:x_{i}>0\right\}$. Each chart is mapped to $\mathbb{R}^{4}$ via $\psi_{i}(x_{1},x_{2},x_{3},x_{4},x_{5})=\left(\frac{x_{j}}{x_{i}},\frac{x_{k}}{x_{i}},\frac{x_{l}}{x_{i}},\frac{x_{m}}{x_{i}}\right)$, where the indices $j$, $k$, $l$, and $m$ label four mutually different coordinates, ordered as $j<k<l<m$, with the additional requirement that none of them coincides with the chart index $i$.
On the upper four sphere $\mathbb{S}^{4+}$, the hypersurface $x^{2}+y^{2}=0$ is mapped to the condition $x_{1}^{2}=x_{2}^{2}$, which, together with the spherical constraint, leads to $2x_{1}^{2}+x_{3}^{2}+x_{4}^{2}+x_{5}^{2}=1$. Likewise, the hypersurface $x^{2}+y^{2}=1$ transforms into $x_{1}^{2}+x_{2}^{2}=x_{5}^{2}$, leading to $2x_{2}^{2}+x_{3}^{2}+x_{4}^{2}=1$, together with $0\leq x_{3}\leq x_{5}\leq 1$ ($0\leq z\leq 1$). The critical points at infinity are located on the equator of the four-sphere, characterized by $x_{5}=0$. Consequently, for the phase-space region under consideration, these points lie on the equator and must satisfy $\Big\{(x_{1},x_{2},x_{3},x_{4},x_{5})\in\mathbb{S}^{4}:x_{1}^{2}=x_{2}^{2},x_{2}\geq 0,x_{3}=0,x_{5}=0\Big\}$.
Focusing on the first local chart, we introduce the coordinates $(\chi,\zeta,\xi,\eta)$ through the mapping $\psi_{1}(x_{1},x_{2},x_{3},x_{4},x_{5})=(\chi,\zeta,\xi,\eta)$. In terms of these variables, the autonomous system (24) is correspondingly transformed into a new dynamical system
$\chi_{\mu}$ $=\frac{1}{2}\chi\eta\Bigl\{6\eta^{3}+\sqrt{6}\eta\xi(\chi^{2}-1)+2(\alpha-3)\eta^{2}\zeta$
$-\zeta(\chi^{2}-1)(\sqrt{6}\xi+2\alpha)\Bigr\},$ (53a)
$\zeta_{\mu}$ $=\frac{1}{2}\zeta\Bigl\{6\eta^{4}+\eta^{2}\left[3(2+\zeta^{2})+\chi^{2}(6+\sqrt{6}\xi)\right]$
$+\zeta^{2}(3+3\chi^{2}-4\alpha)+(2\alpha-9)\eta^{3}\zeta$
$-\eta\zeta\left[9-6\alpha+\chi^{2}(9+2\alpha+\sqrt{6}\xi)\right]\Bigr\},$ (53b)
$\xi_{\mu}$ $=\frac{1}{2}\eta\xi\Bigl\{3\eta^{3}+\eta\left[3+\chi^{2}(3+\sqrt{6}\xi)\right]+(2\alpha-3)\eta^{2}\zeta$
$-\zeta\left[3-2\alpha+\chi^{2}(3+2\alpha+\sqrt{6}\xi)\right]$
$+2\sqrt{6}\delta(\eta-\zeta)(\overline{\lambda}\eta-\xi)\Bigr\},$ (53c)
$\eta_{\mu}$ $=\frac{1}{2}\eta^{2}\Bigl\{3\eta^{3}+\eta\left[3+\chi^{2}(3+\sqrt{6}\xi)\right]+(2\alpha-3)\eta^{2}\zeta$
$-\zeta\left[3-2\alpha+\chi^{2}(3+2\alpha+\sqrt{6}\xi)\right]\Bigr\},$ (53d)
where the “time” variable $\tau$ was rescaled, $d\tau=\eta^{4}d\mu$, in order to remove the divergence arising for $\eta=0$.
To investigate the structure of the phase space at infinity, we set $\eta=0$. The dynamical system (53) reduces to $\chi_{\tau}=\xi_{\tau}=\eta_{\tau}=0$ and $\zeta_{\tau}=\zeta^{3}(3-4\alpha+3\chi^{2})/2$, yielding, for our phase space region, a continuous family of critical points $(1,0,\xi,0)$, with $\xi\in\mathbb{R}$. These points correspond, on the compactified phase space $\mathbb{S}^{4}$, to $A^{\infty}\Bigl\{\left(x_{1},x_{2},x_{3},x_{4},x_{5}\right)\in\mathbb{S}^{4}:x_{1}^{2}=x_{2}^{2},x_{1}>0,x_{2}>0,x_{3}=0,x_{5}=0\Bigr\}$. As all eigenvalues vanish at $(1,0,\xi,0)$, a linear analysis fails to determine stability. Nevertheless, restricting the flow to the invariant manifold $\zeta=0$, one finds $\chi_{\tau}=3\eta^{4}\chi+\sqrt{6}\eta^{2}\chi\xi(\chi^{2}-1)/2$, which reduces to $\chi_{\tau}\simeq 3\eta^{4}$ in the vicinity of $\chi=1$. This implies that $\chi$ always increases, leading to the conclusion that the set $A^{\infty}$ is unstable.
In the second chart, we adopt the local coordinates $(\chi,\zeta,\xi,\eta)$ defined by the map $\psi_{2}:(x_{1},x_{2},x_{3},x_{4},x_{5})\rightarrow(\chi,\zeta,\xi,\eta)$. With this choice, the autonomous system (24) becomes
$\chi_{\mu}$ $=-\frac{1}{2}\eta\Bigl\{6\eta^{3}\chi-\sqrt{6}\eta\xi(\chi^{2}-1)+2(\alpha-3)\eta^{2}\chi\zeta$
$+\zeta(\chi^{2}-1)(\sqrt{6}\xi+2\alpha\chi)\Bigr\},$ (54a)
$\zeta_{\mu}$ $=-\frac{1}{2}\zeta(\eta-\zeta)\Bigl\{3\eta^{2}\zeta-\eta(6+6\chi^{2}+\sqrt{6}\chi\xi)$
$+\zeta\left[3+(3-4\alpha)\chi^{2})\right]\Bigr\},$ (54b)
$\xi_{\mu}$ $=-\frac{1}{2}\eta\xi(\eta-\zeta)\Bigl\{3\eta^{2}-3-3\chi^{2}+\sqrt{6}(2\delta-1)\chi\xi$
$-2\sqrt{6}\delta\overline{\lambda}\eta\chi\Bigr\},$ (54c)
$\eta_{\mu}$ $=-\frac{1}{2}\eta^{2}(\eta-\zeta)(3\eta^{2}-3-3\chi^{2}-\sqrt{6}\chi\xi).$ (54d)
For $\eta=0$, this system simplifies to $\chi_{\tau}=\xi_{\tau}=\eta_{\tau}=0$ and $\zeta_{\tau}=\zeta^{3}\left[(3-4\alpha)\chi^{2}+3\right]/2$, giving rise, within our phase space region, to the critical point $(\pm 1,0,\xi,0)$, with $\xi\in\mathbb{R}$.
On the compactified phase space $\mathbb{S}^{4}$, these points correspond to $B^{\infty}\Bigl\{\left(x_{1},x_{2},x_{3},x_{4},x_{5}\right)\in\mathbb{S}^{4}:x_{1}^{2}=x_{2}^{2},x_{2}>0,x_{3}=0,x_{5}=0\Bigr\}$. Since the linearized system has only zero eigenvalues, stability cannot be inferred from linear theory. However, along the invariant surface $\zeta=0$, the evolution of $\chi$ near $\chi=\pm 1$ is governed by $\chi_{\tau}=\mp 3\eta^{4}$, which confirms the unstable character of $B^{\infty}$. In addition, one can also check that $A^{\infty}\subset B^{\infty}$.
In the third local chart, the constraint $x_{3}>0$ is imposed. However, the critical points at infinity are required to satisfy $\Big\{(x_{1},x_{2},x_{3},x_{4},x_{5})\in\mathbb{S}^{4}:x_{1}^{2}=x_{2}^{2},x_{2}\geq 0,x_{3}=0,x_{5}=0\Big\}$. Since this set is disjoint from the region $x_{3}>0$, no critical points at infinity exist in the third chart.
For the fourth chart, we adopt the map $\psi_{4}(x_{1},x_{2},x_{3},x_{4},x_{5})=(\chi,\zeta,\xi,\eta)$. The autonomous system (24) assumes the equivalent representation
$\chi_{\mu}$ $=\frac{1}{2}\eta\Bigl\{-\eta(3\chi^{3}+\sqrt{6}\zeta^{2}+3\chi\zeta^{2})+(3-2\alpha)\eta^{2}\chi\xi$
$-3\eta^{3}\chi+\xi\left[\sqrt{6}\zeta^{2}+(3-2\alpha)\chi^{3}+(3+2\alpha)\chi\zeta^{2}\right]$
$-2\sqrt{6}\delta\chi^{2}(\eta-\xi)(\overline{\lambda}\eta-1)\Bigr\},$ (55a)
$\zeta_{\mu}$ $=\frac{1}{2}\eta\zeta(\eta-\xi)\Bigl\{3\eta^{2}-3\chi^{2}-3\zeta^{2}+\sqrt{6}(2\delta-1)\chi$
$-2\sqrt{6}\delta\overline{\lambda}\eta\chi\Bigr\},$ (55b)
$\xi_{\mu}$ $=\frac{1}{2}\xi(\eta-\xi)\Bigl\{3\eta^{3}+3\eta(\chi^{2}+\zeta^{2})-3\eta^{2}\xi-3\zeta^{2}\xi$
$+(4\alpha-3)\chi^{2}\xi-2\sqrt{6}\delta\eta\chi(\overline{\lambda}\eta-1)\Bigr\},$ (55c)
$\eta_{\mu}$ $=-\sqrt{6}\delta\eta^{2}\chi(\eta-\xi)(\overline{\lambda}\eta-1).$ (55d)
Setting $\eta=0$, the dynamical system reduces to $\chi_{\tau}=\zeta_{\tau}=\eta_{\tau}=0$ and $\xi_{\tau}=\xi^{3}\left[3\zeta^{2}+(3-4\alpha)\chi^{2}\right]/2$. This yields a family of relevant critical points at infinity for the considered phase-space region, given by $(\chi,\zeta,0,0)$ with $\chi^{2}=\zeta^{2}$. On the compactified phase space $\mathbb{S}^{4}$, these points correspond to $\Bigl\{\left(x_{1},x_{2},x_{3},x_{4},x_{5}\right)\in\mathbb{S}^{4}:x_{1}^{2}=x_{2}^{2},x_{2}\geq 0,x_{3}=0,x_{4}>0,x_{5}=0\Bigr\}$, which can be decomposed into $\Bigl\{\left(x_{1},x_{2},x_{3},x_{4},x_{5}\right)\in\mathbb{S}^{4}:x_{1}^{2}=x_{2}^{2},x_{2}>0,x_{3}=0,x_{4}>0,x_{5}=0\Bigr\}\cup(0,0,0,1,0)$. It is straightforward to verify that the first subset is contained in $B^{\infty}$. Consequently, this chart admits an additional critical point at infinity, namely $C^{\infty}(0,0,0,1,0)$. Since the linearized system has only zero eigenvalues, stability cannot be inferred from linear theory. However, restricting the flow to the invariant manifold $\xi=0$, one finds $\chi_{\tau}=-\sqrt{6}\zeta^{2}\eta^{2}/2$ in the vicinity of $\chi=0$. Since $\chi$ always decreases, the point $C^{\infty}$ is therefore unstable.
The analysis in the fifth chart reproduces a dynamical structure analogous to that of the system (24), without introducing any further critical points at infinity.
Overall, our analysis demonstrates that the autonomous system (24) contains only two critical points at infinity, $B^{\infty}$ and $C^{\infty}$, both displaying unstable behavior. Therefore, we conclude that the coupled phantom cosmological model with an interaction term $Q$ given by Eq. (20), with a constant dissipation coefficient, effectively avoids Big-Rip singularities.
In the case of a variable dissipation coefficient, a similar analysis in all respects reveals the existence of the same critical points, with the same stability properties, leading, consequently, to the same conclusion.